suppose that P(x)=x^2-5x+2.
Find P(0) and P(3)

Answer by Theo(13342)

suppose that P(x)=x^2-5x+2.
Find P(0) and P(3)
to find p(0), replace x with 0 to get:

(0)^2 - 5*(0) + 2 = 0 + 0 + 2 = 2

to find p(3), replace x with 3 to get:

(3)^2 - 5*(3) + 2 = 9 - 15 + 2 = -4